# Evolving Large Language Model Assistant with Long-Term Conditional Memory

### v1

@@ Abstract
With the rapid development of large language models, AI assistants like ChatGPT have widely entered people’s works and lives. In this paper, we present an evolving large language model assistant that utilizes verbal long-term memory. It focuses on preserving the knowledge and experience from the history dialogue between the user and AI assistant, which can be applied to future dialogue for generating a better response. The model generates a set of records for each finished dialogue and stores them in the memory. In later usage, given a new user input, the model uses it to retrieve its related memory to improve the quality of the response. To find the best form of memory, we explore different ways of constructing the memory and propose a new memorizing mechanism called conditional memory to solve the problems in previous methods. We also investigate the retrieval and usage of memory in the generation process. The assistant uses GPT-4 as the backbone and we evaluate it on three constructed test datasets focusing on different abilities required by an AI assistant with long-term memory.
Introduction
@@
Unfortunately, one main problem for the current AI assistant is that the information in one dialogue only last for one dialogue session. When a user starts a new session, the AI assistant is reset to its initial state and it will not preserve any information from previous dialogue sessions. This prevents the AI assistant from acquiring knowledge and experience from the long-term dialogue, to improve the quality of its responses, like humans do. As shown in the example in Figure 1, when the user asks the assistant to help him to capitalize some words, the assistant chooses to capitalize the first letter of each word, which is not what the user expected. Then the user provides more detailed feedback to avoid ambiguity and get the answer he wants. If the assistant is a human, he can learn that “capitalize” refers to capitalizing each letter for this user and can give the right response in the later dialogue, while a LLM assistant fails to achieve this. Therefore, our aim is to explore how to endow the LLM assistant with such an essential capability that allows it to learn knowledge and experience from previous dialogues and apply them to later ones.
@@
Figure 1: An example about the evolving LLM assistant.
@@
In this work, we propose to adopt a memory-based framework as a wrapper outside the LLM assistant to address this challenge. The evolving LLM assistant mainly consists of three parts: an existing LLM assistant, a memory, and a prompt-based wrapper conducting interactions between the assistant and the memory. As shown in Figure 2, the wrapper utilizes the ongoing dialogue and LLM assistant itself to construct memory records and store them in the memory. Every time the assistant generates a response, we use a retrieval model to retrieve a series of memory records related to the user input from the memory and combine the information and the user input to generate a better response. Using a memory-based framework has a few advantages. This method does not require additional finetune of the LLM, which is crucial when most of the current LLMs are unattainable or expensive to finetune. It also allows the framework to be easily deployed on any device with an LLM API, without the need for a GPU. Moreover, the memory file is the only place that stores user-specific information, so it is private and can be easily transferred between users’ own devices.
@@
To further explore the details of the memory-based evolving LLM assistant, we conduct our research from three aspects. Firstly, the memorizing mechanism is a key point for a memory-based model. The way to convert a large amount of dialogue history into effective memory records is crucial for the performance of the entire model. Two straightforward ideas based on history and summaries from previous works (Xu, Szlam, and Weston 2021; Park et al. 2023) are tested under this setting. To further address the shortcomings of these two intuitive ideas, we propose conditional memory, where we believe that the memorization timing and the content of the memory records should be conditional. We also present a self-reflection mechanism for memory retrieval, which allows the LLM assistant to determine whether the retrieved information is sufficient to respond and whether to conduct further searches. Considering there is no existing dataset that is appropriate for testing the ability of an LLM assistant with long-term memory, we build three small-scale testing datasets using GPT-4 from different ability aspects: continuing previous dialogue, learning new knowledge, and learning from human feedback. Our contribution can be concluded as follow:
  • We propose a memory-based dialogue system as a solution for an evolving LLM assistant that continually improves itself while the user uses it.
  • We investigate the framework of the evolving LLM assistant and propose a new memorizing mechanism.
  • We construct three test datasets on different aspects to test the model’s abilities in learning from the dialogue history.
Figure 2: The framework of the evolving LLM assistant.
Related Work
@@
Retrieval-based dialogue system is a popular topic in the field and has long been studied as an important paradigm. Open-domain chit-chat dialogue system is a typical embodiment of this paradigm. Such dialogue systems are challenging due to the diversity of possible responses to a single dialogue history. Hence, external knowledge is necessary for avoiding safe but boring responses. These dialogue systems usually retrieve information from the dialogue corpora and focus on how to integrate the retrieved dialogue into the generation process (Wu et al. 2019; Cai et al. 2018; Yang et al. 2019; Gupta et al. 2020; Paranjape et al. 2021). Other work further extends the external information to different types of knowledge such as knowledge bases and external documents (Dinan et al. 2018; Lian et al. 2019; Qin et al. 2019; Wu et al. 2021; Komeili, Shuster, and Weston 2021).
@@
Another type of work using the paradigm is conversational question answering. Considering it is a combination of both dialogue and QA, a retrieval system that searches the answer information from the knowledge base or document corpus is inevitable for improving the accuracy of the answer. Different from the open-domain dialogue system, conversational QA pays more attention to the accuracy of the retrieval including the retrieval optimization and query understanding (Qu et al. 2019; Christmann et al. 2019; Qu et al. 2020; Vakulenko et al. 2021; Adlakha et al. 2022).
@@
Recently, as a stronger version of the dialogue system, LLM-based chatbots such as ChatGPT can also be augmented by retrieval. These systems are capable of providing a more informative and faithful answer by retrieving question-related web documents directly from the Internet (He, Zhang, and Roth 2022; Ram et al. 2023; Liu et al. 2023). There also exist many toolkits for building this type of system such as LangChain.
@@
Memory LLM Call Length
Example
History 1 26.3
Assistant: Do you prefer going to the gym? User: I like the controlled environment. I don’t want to have to depend on the weather considering where I live.
Summary 1.2 74.6
The user and the assistant discussed their preferences for running and going to the gym. … They also discussed the availability of parks and running trails in the area and the possibility of joining or starting a running club.
Conditional 1.5 46.8
Context: The user and the assistant are discussing the inconvenience of running outdoors due to weather conditions. Knowledge: User likes gym workouts due to a controlled environment and not having to depend on weather.
Table 1: Statistics and examples of the three types of memory generated from chit-chat dialogue in multi-session chat dataset (Xu, Szlam, and Weston 2021). “LLM Call” refers to the average LLM api call per utterance (response and memory construction) for the LLM assistant, and “Length” represents the average length of the generated memory records.
Memory-based models can be considered a special type of retrieval-based model. Instead of retrieving information from an existing source, the memory-based models retrieve useful information from its memory while keeping add more memory records to it. This also makes how to construct the memory a key point in the research of this type of model. Some researchers (Park et al. 2023) apply the memory-based dialogue system on interactive agents where some AI agents simulate humans in a virtual environment. A history-based memory system is used to help the agents achieve long-term interaction. Other researchers (Xu, Szlam, and Weston 2021) use it on continuing a multi-session dialogue and propose a summary-based memory system. In our work, we also adopt a solution based on a memory system and target its application in building an evolving LLM assistant.
Method
The memory-based framework of the evolving LLM assistant consists of two parts: the memory construction during the dialogue and the retrieval and use of the memory records in response generation. In this section, we mainly focus on the methods and more details such as prompt design are shown in the Appendix.
Memory Construction
The construction of memory is a crucial problem in a memory-based model. The varying forms or sources of memory can significantly affect the efficacy of the model. In our research, we investigate three distinct memory types and compare their respective influence on the model’s performance.
@@
Dialogue summaries can also be used as memory. As stated by (Xu, Szlam, and Weston 2021), using dialogue summaries as memory provides more context information than using dialogue history for the retrieval and generation process. In this work, we use GPT-4 to generate a summary for each dialogue session. Here, a dialogue session stands for multiple utterances that are continuous in time and content. These generated summaries are stored as memory records.
@@
Based on our assumption that dialogue information is conditional, we propose to divide one memory record into two parts: context and knowledge. Context is a summary that describes the background of the target utterance based on its neighbors, and knowledge is the high-level conclusion about what the assistant learns from the target utterance. By setting different objectives for the two parts, we can achieve a more comprehensive memory record that takes into account both retrieval and generation abilities. We also use the GPT-4 to generate the two parts with different instruction prompts.
@@
There is no limitation that only one type of memory record is allowed in the memory. It is possible that a single type of memory record is not sufficient to effectively contain all information related to user input. In this case, we also try to find out whether the multi-view memory provides more diverse information for the response generation.
Memory Retrieval and Application
@@
We use dense retrieval to search the related memory records based on the user input. We encode the user input and each memory record to a vector using SimCSE model (Gao, Yao, and Chen 2021). Considering the scale of the memory is not large enough to use the FAISS, we simply calculate the cosine similarity between the representation of the memory record and the user input and use the top $K$ records for response generation.
@@
Dataset Num Sour Sour Test
@@
Table 2: The statistics of the three test datasets. “Session Sour” refers to the average session number of the source dialogue, “Turn Sour” represents the average utterance number from all source sessions and “Turn Test” represents the average utterance number for testing.
@@
To evaluate an LLM assistant with memory, we construct test datasets from three different aspects: continuing previous dialogue, learning new knowledge, and learning from user feedback. Data samples from the three datasets follow the same structure, which contains a dialogue history as the source for constructing memory and some testing dialogue utterance with reference responses. Due to the high cost of GPT-4 usage, the three datasets are only small-scale datasets for testing. The statistics of the datasets are shown in Table 2. Here, we briefly introduce the three datasets and their construction process. More details are listed in the appendix.
Continuing Previous Dialogue
@@
For a dialogue system with memory, continuing a previous dialogue session is considered a basic ability. It requires the model to use the information from the previous dialogue to enhance the diversity or readability of the next response. For example, in Figure 4, when a question related to heatmap is asked by the user, the assistant can utilize the project discussed under the topic of heatmap before as an example in its response. To construct such a dataset, we convert it to a dialogue continue writing problem. Given the dialogue between the user and the AI assistant, we use the GPT-4 to continue writing a new utterance that contains information from the existing context. This generated utterance is used for testing and the existing dialogue is used as the history for memory construction. In practice, we adopt a code discussion dialogue between AI assistant and user (Yang et al. 2023) as source dialogue, since it has rich contextual connections.
Learning New Knowledge
@@
Another important ability of the evolving assistant is acquiring unknown knowledge from its dialogue with the user. Knowledge usually has two types. The first type is the unknown/wrong factual knowledge for the LLM. However, for these knowledge, directly retrieving information from the Internet seems a better solution. The second type is the personal knowledge related to the user himself, which can only be learned within the dialogue. Hence, we mainly focus on the ability of the assistant to obtain the user’s personal knowledge. We construct such data based on a multi-session chat (MSC) dataset (Xu, Szlam, and Weston 2021). The MSC dataset contains a multi-session chit-chat between two human and their personal settings. And these personal settings can be inferred from the dialogue. In our dataset, as shown in Figure 4, we assume that one of the two participants in the dialogue is the user and the other one is the assistant. These dialogue sessions are considered dialogue history. Then we use GPT-4 to convert the personal settings of the user to question-answer utterances aiming to test whether the assistant learns these knowledge in the dialogue.
Dataset Conditional+History vs Conditional Conditional+Summary vs Conditional
@@
Table 3: The evaluation result of the multi-view memory. For evaluation metrics, we adopt GPT-score for learning new knowledge, and the Good/Same/Bad test (GSB) for learning from feedback and continuing previous dialogue.
Learning from Human Feedback
The last ability we test is whether the evolving assistant can learn from the user’s feedback and use it properly in a similar situation. For instance, in Figure 4, the user seeks for information about table tennis and provides feedback about requiring more information on how to play it. In the future, when the user asks for information about another sport football, the assistant can directly add information about how to play it in its response. In terms of dataset construction, given a seed instruction/question sampled from the seed set of self-instruction (Wang et al. 2022), we first generate a response using GPT-4 and generate suggestions to improve the response. These suggestions are regarded as pseudo human feedback. Then we use the suggestions to improve the response. In this case, we obtain a two-turn dialogue with human feedback. We also add a background dialogue before it with GPT-4 using a similar method with (Yang et al. 2023). The concatenation of the two dialogues is used as dialogue history. For testing, we generate a similar instruction or question with the seed one and generate a reference response that also follows suggestions given in dialogue history. The assistant is expected to learn from previous feedback and directly generate the reference response in the later dialogue.
Experiment
Application Details
To make our result and evaluation consistent across all the experiments, we adopt the GPT-4 as the backbone of the LLM assistant and dataset construction. For the evaluation metrics, we use an N-gram similarity metric, Rouge (Lin 2004) including Rouge-1 and Rouge-2. We also use the GPT-4 for more complicated evaluations, which are shown in the next subsection. For the hyperparameter, we set the number of the retrieved memory records $K$ as 3 for history-based memory and conditional memory, and we set $K$ as 1 for summary-based memory. It is also worth to notice some source dialogue history can be relatively short. To simulate the long-term use of an assistant, we merge the memory of every 10 data samples into one in the experiment of learning from feedback and continuing previous dialogue.
Automated Evaluation with GPT-4
There are three different types of GPT-4 evaluation in our experiment: scoring, comparing, and multiple choice. The detailed prompt for evaluation is listed in the appendix.
Scoring focuses on giving a predicted response a score between a predefined margin when the reference answer is clear. We call it GPT-score and use it to evaluate the model ability in learning new knowledge. The GPT-4 is required to give a 0-2 score for a response. Here, “0” represents the response is totally wrong, “1” represents the response contains part of personal information in the reference and “2” represents the response contains all information.
Comparison aims to conduct a Good/Same/Bad test (GSB) between two candidates, which is applied in a more complicated evaluation situation. The GPT-4 is used to make a decision about which candidate is better given the reference response and other necessary information. Note that the GPT-4 has a position bias and candidates at a specific position (first or second) may have more advantages. Hence, we conduct the evaluation twice and exchange the position of the two candidates. Only when the two evaluation meets the result of each other, the result is definite. If the results are different, we consider it as a tie.
Multiple choice is similar to comparison, but the GPT-4 needs to choose the best candidate among multiple options. For the position bias problem, we solve it by randomly placing all candidates in different positions.
Experiment Results

### v2

# Personalized Large Language Model Assistant with Evolving Conditional Memory

@@ Abstract
With the rapid development of large language models, AI assistants like ChatGPT have become increasingly integrated into people’s works and lives but are limited in personalized services. In this paper, we present a plug-and-play framework that could facilitate personalized large language model assistants with evolving conditional memory. The personalized assistant focuses on intelligently preserving the knowledge and experience from the history dialogue with the user, which can be applied to future tailored responses that better align with the user’s preferences. Generally, the assistant generates a set of records from the dialogue dialogue, stores them in a memory bank, and retrieves related memory to improve the quality of the response. For the crucial memory design, we explore different ways of constructing the memory and propose a new memorizing mechanism named conditional memory. We also investigate the retrieval and usage of memory in the generation process. We build the first benchmark to evaluate personalized assistants’ ability from three aspects. The experimental results illustrate the effectiveness of our method.
Personalized Large Language Model Assistant with Evolving Conditional Memory
Ruifeng Yuan, Shichao Sun, Yongqi Li, Zili Wang, Ziqiang Cao, Wenjie Li The Hong Kong Polytechnic University INF Technology (Shanghai) co., Ltd. Institute of Artificial Intelligence, Soochow University, China
1 Introduction
@@
Unfortunately, one main problem for the current AI assistant is that the information in one dialogue only lasts for one dialogue session. When a user starts a new session, the AI assistant is reset to its initial state and will not preserve any information from previous dialogue sessions. This prevents the AI assistant from acquiring knowledge and experience from the long-term dialogue, to improve the quality of its responses, like humans do. As shown in the example in Figure 1, when the user asks the assistant to help him capitalize some words, the assistant chooses to capitalize the first letter of each word, which is not what the user expected. Then the user provides more detailed feedback to avoid ambiguity and get the answer he wants. If the assistant is a human, he can learn that “capitalize" refers to capitalizing each letter according to this user’s preference and can give the more suitable response in the later dialogue, while a LLM assistant fails to achieve this. Therefore, our aim is to explore how to endow the LLM assistant with such an essential capability that allows it to learn knowledge and experience from previous dialogues and apply them to later ones.
@@
Figure 1: An example of what a personalized LLM assistant with evolving memory can do.
@@
In this work, we propose to adopt a memory-based framework as a wrapper outside the LLM assistant to address this challenge. Such personalized LLM assistant mainly consists of three parts: an existing LLM assistant, an evolving memory, and a prompt-based framework conducting interactions between the assistant and the memory. As shown in Figure 2, the framework utilizes the ongoing dialogue and LLM assistant itself to construct memory records and store them in the memory. Every time the assistant generates a response, we use a retrieval model to retrieve a series of memory records related to the user input from the memory and combine the information with the user input to generate a better response. Using a memory-based framework has a few advantages. This method does not require additional finetune of the LLM, which is crucial when most of the current LLMs are unattainable or expensive to finetune. It also allows the framework to be easily deployed on any device with an LLM API, without the need for a GPU. Moreover, the memory file is the only place that stores user-specific information, so it is private and can be easily transferred between users’ own devices.
@@
To further explore the details of the memory-based personalized LLM assistant, we conduct our research from three aspects. 1) The memorizing mechanism is a key point for a memory-based model. The way to convert a large amount of dialogue history into effective memory records is crucial but challenging for the performance of the entire model. Two straightforward ideas based on history and summaries from previous works Xu et al. (2021); Park et al. (2023) are tested under this setting. By combining the advantages of both intuitive ideas, we propose conditional memory, where we believe that the memorization timing and the content of the memory records should be conditional. This allows the memory to capture detailed information while being relatively concise. 2) We also present a self-reflection mechanism for memory retrieval, which allows the LLM assistant to determine whether the retrieved information is sufficient to respond and whether to conduct further searches. 3) Considering there is no existing dataset that is appropriate for testing the ability of an LLM assistant with long-term memory, we build three datasets using GPT-4 with different ability aspects: continuing previous dialogue, learning new knowledge, and learning from human feedback.
Our contribution can be concluded as follows: (1) We propose a plug-and-play framework to facilitate the personalized LLM assistant with external evolving memory. (2) For the crucial memory design in personalized LLM assistants, we propose an effective conditional memory mechanism and a self-reflection memory retrieval strategy. (3) We construct the first benchmark to evaluate the personalized assistants’ ability to capture user preferences for tailored responses.
Figure 2: The framework of personalized assistant with evolving memory.
2 Related Work
@@
Retrieval-based dialogue system is a popular topic in the field and has long been studied as an important paradigm. Open-domain chit-chat dialogue system is a typical embodiment of this paradigm. Such dialogue systems are challenging due to the diversity of possible responses to a single dialogue history. Hence, external knowledge is necessary for avoiding safe but boring responses. These dialogue systems usually retrieve information from the dialogue corpora and focus on how to integrate the retrieved dialogue into the generation process Wu et al. (2019); Cai et al. (2018); Yang et al. (2019); Gupta et al. (2020); Paranjape et al. (2021). Other work further extends the external information to different types of knowledge such as knowledge bases and external documents Dinan et al. (2018); Lian et al. (2019); Qin et al. (2019); Wu et al. (2021); Komeili et al. (2021).
@@
Another type of work using the paradigm is conversational question answering. Considering it is a combination of both dialogue and QA, a retrieval system that searches the answer information from the knowledge base or document corpus is inevitable for improving the accuracy of the answer. Different from the open-domain dialogue system, conversational QA pays more attention to the accuracy of the retrieval including the retrieval optimization and query understanding Qu et al. (2019); Christmann et al. (2019); Qu et al. (2020); Vakulenko et al. (2021); Adlakha et al. (2022).
@@
Recently, as a stronger version of the dialogue system, LLM-based chatbots such as ChatGPT can also be augmented by retrieval. These systems are capable of providing a more informative and faithful answer by retrieving question-related web documents directly from the Internet He et al. (2022); Ram et al. (2023); Liu et al. (2023a). There also exist many toolkits for building this type of system such as LangChain.
@@
Memory Frequency Length Example
History 1 52.6 Assistant: Do you prefer going to the gym? User: I like the controlled environment. I don’t want to depend on the weather.
Summary 0.2 74.6 The user and the assistant discussed their preferences for running and going to the gym. … They also discussed the availability of parks and running trails in the area.
Conditional 0.5 46.8 Context: The user and the assistant are discussing the inconvenience of running outdoors due to weather conditions. Knowledge: User likes gym workouts due to a controlled environment and not having to depend on weather.
Table 1: Statistics and examples of the three types of memory generated from chit-chat dialogue in multi-session chat dataset Xu et al. (2021). “Frequency" refers to the average number of memory records generated per utterance. “Length" represents the average length of the generated memory records.
Memory-based models can be considered a special type of retrieval-based model. Instead of retrieving information from an existing source, the memory-based models retrieve useful information from its memory while keeping add more memory records to it. This also makes how to construct the memory a key point in the research of this type of model. Some researchers Park et al. (2023) apply the memory-based dialogue system on interactive agents where some AI agents simulate humans in a virtual environment. A history-based memory system is used to help the agents achieve long-term interaction. Other researchers Xu et al. (2021) use it on continuing a multi-session dialogue and propose a summary-based memory system. In our work, we also adopt a solution based on a memory system and target its application in building an personalized LLM assistant.
3 Method
The memory-based framework of the personalized LLM assistant consists of two parts: the memory construction during the dialogue and the retrieval and use of the memory records in response generation. In this section, we mainly focus on the methods and more details such as prompt design are shown in the Appendix.
3.1 Memory Construction
The construction of memory is a crucial problem in a memory-based model. The varying forms or sources of memory can significantly affect the efficacy of the model. In our research, we investigate three distinct memory types and compare their respective influence on the assistance’s performance. The example and statistic of the three types of memory is shown in Table 1.
@@ 3.1 Memory Construction
Dialogue summaries can also be used as memory. As stated by Xu et al. (2021), using dialogue summaries as memory provides more context information than using dialogue history for the retrieval and generation process. In this work, we use GPT-4 to generate a summary for each dialogue session. Here, a dialogue session stands for multiple utterances that are continuous in time and content. These generated summaries are stored as memory records.
@@
Based on our assumption that dialogue information is conditional, we propose to divide one memory record into two parts: context and knowledge. Context is a summary that describes the background of the target utterance based on its neighbors, and knowledge is the high-level conclusion about what the assistant learns from the target utterance. By setting different objectives for the two parts, we can achieve a more comprehensive memory record that takes into account both retrieval and generation abilities. We use the GPT-4 to generate the two parts with different prompts.
@@
There is no limitation that only one type of memory record is allowed in the memory. It is possible that a single type of memory record is not sufficient to contain all information related to user input. Here, we also try to find out whether the multi-view memory provides more diverse information for the response generation.
3.2 Memory Retrieval and Application
@@
We use dense retrieval to search the related memory records based on the user input. We encode the user input and each memory record to a vector using SimCSE model Gao et al. (2021). Considering the scale of the memory is not large enough to use the FAISS, we calculate the cosine similarity between the representation of the memory record and the user input and use the top $K$ records for response generation.
@@ 4 Dataset
Dataset Num Source Source Test
@@
Table 2: The statistics of datasets. “Session Source" and “Turn Source" refers to the average session number and utterance number of source dialogues.“Turn Test" represents the average utterance number for testing.
@@
To evaluate a personalized assistant, we construct datasets from following aspects: continuing previous dialogue, learning new knowledge, and learning from user feedback. Data samples from the three datasets follow the same structure, which contains a dialogue history as the source for constructing memory and some testing dialogue utterance with reference responses. All the three datasets are synthetic data based on GPT-4. In this case, due to the high cost of GPT-4 usage, the datasets are only small-scale datasets. The statistics of the datasets are shown in Table 2. Here, we briefly introduce the three datasets and their construction process. The human evaluation for datasets and construction details are in the appendix. This allows the researchers to generate more data samples with GPT-4.
4.1 Continuing Previous Dialogue
@@
For a personalized assistant with memory, continuing a previous dialogue session is a basic ability. It requires the model to use information from previous dialogues to enhance the diversity or readability of the next response. For example, in Figure 4, when a question related to heatmap is asked by the user, the assistant can utilize the project discussed under the topic of heatmap before as an example in its response. To construct such a dataset, we convert it to a dialogue continue writing problem. Given the dialogue between the user and the AI assistant, we use the GPT-4 to continue writing a new utterance that contains information from the existing context. This generated utterance is used for testing and the existing dialogue is used as the history for memory construction. In practice, we adopt a code discussion dialogue between AI assistant and user Yang et al. (2023) as source dialogue, since it has rich contextual connections.
4.2 Learning New Knowledge
@@
Another important ability of the personalized assistant is acquiring personal knowledge from its dialogue with the user. We construct such data based on a multi-session chat (MSC) dataset Xu et al. (2021). The MSC dataset contains a multi-session chit-chat between two human and their personal settings. And these personal settings can be inferred from the dialogue. In our dataset, as shown in Figure 4, we assume that one of the two participants in the dialogue is the user and the other one is the assistant. These dialogue sessions are considered dialogue history. Then we use GPT-4 to convert the personal settings of the user to question-answer utterances aiming to test whether the assistant learns these knowledge in the dialogue.
4.3 Learning from Human Feedback
The last ability we test is whether the personalized assistant can learn from the user’s feedback or preference and use it properly in a similar situation. For instance, in Figure 4, the user seeks for information about table tennis and provides feedback about requiring more information on how to play it. In the future, when the user asks for information about another sport football, the assistant can directly add information about how to play it in its response. In terms of dataset construction, given a seed instruction/question sampled from the seed set of self-instruction Wang et al. (2022), we first generate a response using GPT-4 and generate suggestions to improve the response. These suggestions are regarded as pseudo human feedback. Then we use the suggestions to improve the response. In this case, we obtain a two-turn dialogue with human feedback. We also add a background dialogue before it with GPT-4 using a similar method with Yang et al. (2023). The concatenation of the two dialogues is used as dialogue history. For testing, we generate a similar instruction or question with the seed one and generate a reference response that also follows suggestions given in dialogue history. The assistant is expected to learn from previous feedback and directly generate the reference response in the later dialogue.
5 Experiment
5.1 Application Details
Considering we aim to test the effectiveness of different memory mechanisms without the limitation of the LLM, we use GPT-4, current strongest LLM, as the backbones of LLM assistants in all experiment. For the evaluation metrics, we use an N-gram similarity metric, Rouge Lin (2004) including Rouge-1 and Rouge-2. We also use the GPT-4 for more complicated evaluations, which are shown in the next subsection. For the hyperparameter, we set the number of the retrieved memory records $K$ as 3 for history-based memory and conditional memory, and we set $K$ as 1 for summary-based memory. It is also worth to notice some source dialogue history can be relatively short. To simulate the long-term use of an assistant, we merge the memory of every 10 data samples into one in the experiment of learning from feedback and continuing previous dialogue.
Dataset Cond+His vs Cond Cond+Sum vs Cond
@@ 5.1 Application Details
Table 3: The evaluation result of combining conditional memory (Cond) with history (His) and summary (Sum). For evaluation metrics, we adopt GPT-score for learning new knowledge, and GSB test for the other two datasets.
5.2 Automated Evaluation with GPT-4
There are three different types of GPT-4 evaluation in our experiment: scoring, comparing, and multiple choice. To ensure the fairness in GPT-4 evaluation Wang et al. (2023), we list the details and related prompts in the appendix.
Scoring Liu et al. (2023b); Fu et al. (2023) focuses on giving a predicted response a score between a predefined margin when the reference answer is clear. We call it GPT-score. The GPT-4 is required to give a 0-2 score for a response. Here, “0” represents the response is totally wrong, “1” represents the response contains part of information in the reference answer and “2” represents the response contains all information.
Comparison Zheng et al. (2023) aims to conduct a Good/Same/Bad test (GSB) between two candidates, which is applied in a more complicated evaluation situation. The GPT-4 is used to make a decision about which candidate is better given the reference response and other necessary information. Multiple choice is similar to comparison, but the GPT-4 needs to choose the best candidate among multiple options.
We use GPT-score for learning new knowledge. As for the GSB test, it is used for learning from feedback and continuing previous dialogue. The main reason for such difference is that learning new knowledge provides clear reference answers where a direct comparison is more precise, while the other two tasks have more open answers that are more suitable for the GSB test.
5.3 Experiment Results